# Unpacking Positional Encoding in Transformers: A Spectral Analysis of Content-Position Coupling

### v1

@@ 3.1 Preliminaries and Assumptions
It can be verified that all forms of sine-cosine position embedding of absolute position encoding satisfy this assumption and the actual effect of relative position encoding is also presented in the form of a Toeplitz matrix.
Together, these two assumptions enable us to explore how positional information modulates the spectral behavior of attention matrices, and to quantify the effects of different PE strategies using tools from the spectral theory of Toeplitz operators.
3.2 Toeplitz Representation of Positional Encoding
In the Transformer’s attention mechanism, the interaction between representations at positions $i$ and $j$ is captured by the attention logit matrix $\mathbf{L}$, where $L_{i,j}=\text{AttentionLogit}_{i,j}$. Different positional encoding (PE) mechanisms influence this logit matrix in distinct ways to incorporate positional information.
Following our decomposition $x_{i}=c_{i}+p_{i}$, where $c_{i}$ is the content component and $p_{i}$ is the position component, and defining $q_{i}^{c}=W_{Q}c_{i}$, $q_{i}^{p}=W_{Q}p_{i}$, $k_{j}^{c}=W_{K}c_{j}$, and $k_{j}^{p}=W_{K}p_{j}$, we analyze the structure of $\mathbf{L}$.
We denote the Gram matrix of a set of column vectors $\{v_{i}\}$ and $\{w_{j}\}$ as $G_{v,w}$, where $(G_{v,w})_{i,j}=v_{i}^{\top}w_{j}$. Under this notation, for PE methods that operate via additive modification of token representations or attention biases (excluding RoPE), the attention logit matrix can be expressed as:
$\mathbf{L}_{\text{Non-RoPE}}=G_{q^{c},k^{c}}+G_{q^{c},k^{p}}+G_{q^{p},k^{c}}+G_{q^{p},k^{p}}+\mathbf{B}$ (8)
Here, $\mathbf{B}$ denotes an explicit bias matrix (e.g., from T5 or ALiBi). Under our theoretical assumptions, the term $G_{q^{p},k^{p}}$ forms a Toeplitz matrix, since it depends only on relative positions. Similarly, $\mathbf{B}$ is Toeplitz, as it typically encodes relative-position-dependent biases ($B_{i,j}=b_{i-j}$). The remaining terms involving content ($G_{q^{c},k^{c}}$, $G_{q^{c},k^{p}}$, $G_{q^{p},k^{c}}$) are not Toeplitz due to their dependence on specific token content.
In contrast, RoPE applies complex-valued rotations to token representations. By pairing adjacent dimensions into complex numbers, we construct complex-valued vectors $\mathbf{q}_{i}$, $\mathbf{k}_{i}$, and $\mathbf{p}_{i}$, and define a modulation vector $\mathbf{e}_{i}$ such that $\mathbf{e}_{i}^{(j)}=e^{i\theta_{j}}$. The logit matrix then takes the form:
$\mathbf{L}_{\text{RoPE\_complex}}=\left(G_{\mathbf{q},\mathbf{k}}+G_{\mathbf{q},\mathbf{p}}+G_{\mathbf{p},\mathbf{k}}+G_{\mathbf{p}}\right)\circ G_{\mathbf{e}}$ (9)
with $\circ$ denoting the Hadamard (element-wise) product, and the final logit is:
$\mathbf{L}_{\text{RoPE}}=\operatorname{Re}\left\{\mathbf{L}_{\text{RoPE\_complex}}\right\}$ (10)
Since $G_{\mathbf{e}}$ is Toeplitz by construction, the Hadamard product introduces relative-position dependence multiplicatively across all terms.
This formal structure allows us to compare how various PE mechanisms couple content and positional information. In the Non-RoPE case:
  • $G_{q^{c},k^{c}}$ encodes content-content similarity.
  • $G_{q^{p},k^{p}}$ captures position-position interactions and is Toeplitz.
  • $G_{q^{c},k^{p}}$ and $G_{q^{p},k^{c}}$ are cross terms encoding interactions between content and absolute position.
  • $\mathbf{B}$ encodes a content-independent, relative-position bias and is Toeplitz.
While $G_{q^{p},k^{p}}$ and $\mathbf{B}$ directly provide relative-position structure, they do not incorporate content-specific modulation. The cross-terms $G_{q^{c},k^{p}}$ and $G_{q^{p},k^{c}}$ may encode interactions between content and absolute position, but for these to collectively capture content-relative position coupling, the model would need to learn mappings where the cross-term interactions depend solely on $i-j$, yet are modulated by content. Such constrained mappings are difficult to realize in practice.
By contrast, RoPE’s multiplicative structure enables a more direct content-relative interaction: each component in the Gram matrix sum is modulated entrywise by the Toeplitz matrix $G_{\mathbf{e}}$, allowing relative-position structure to directly shape the contribution of content vectors. This results in a more expressive and learnable coupling mechanism between content and relative position.
Based on this analysis, we hypothesize that PE mechanisms enabling direct content-relative interaction within the attention logits will perform better on tasks requiring such coupling. RoPE, via its Hadamard product with $G_{\mathbf{e}}$, provides this mechanism, whereas Non-RoPE methods rely on structurally constrained additive approximations.
To test this hypothesis, we will design a synthetic evaluation task specifically tailored to probe the model’s ability to utilize content-relative positional information. This task will serve as our first benchmark for evaluating the expressive power of different positional encoding strategies.
3.3 Spectral Properties and Theoretical Prediction of Deposit Patterns
The structure of the attention logit matrix $\mathbf{L}$, particularly the properties of its Toeplitz or Toeplitz-like components, provides crucial insights into the stability and dynamics of positional information processing. The spectral properties of these matrices, especially the distribution and range of their eigenvalues, are known to influence the convergence of optimization algorithms.
Theorem 3.1 (Szegő’s Theorem on Eigenvalue Distribution (From grenander2002toeplitz )).
Spectral contraction—manifested as a reduced eigenvalue range or lower condition number—is well known to accelerate and stabilize gradient-based learning boyd2004convex ; gray2006toeplitz . The spectral properties associated with RoPE’s coupling mechanism, through this contraction, facilitate faster and more stable learning of the content-relative positional interaction. We give a more detailed analysis in Appendix B.
This enhanced convergence property leads to a concrete structural prediction: in early layers, the model tends towards specializing attention heads for content-relative positional coupling, while other heads in the same layer may exhibit less specific positional roles. This specialization results in a localized deposit pattern, where certain heads disproportionately contribute to processing positional information. Such behavior is expected to emerge more prominently under PE mechanisms with favorable spectral properties—most notably, RoPE—and to be absent or weaker in less structured encodings.
To validate this prediction, we design ablation experiments to confirm the emergence and characteristics of this deposit pattern.
4 Task Design and Experiment
4.1 Experiment Setup and Task Design
To evaluate our theoretical predictions on positional encoding and learning dynamics, we use a 6-layer Transformer decoder and compare six PE configurations: Absolute Positional Encoding, T5-style Relative Positional Encoding, ALiBi, NoPE (no positional encoding), Rotary Positional Encoding (RoPE), and random, fixed embeddings. Full training details—including hyperparameters, optimizer settings, and vocabulary construction—are provided in Appendix C.
We introduce two synthetic classification tasks that isolate positional requirements:
Task 1: Trigger Word Relative Distance Prediction. Each sequence contains two designated trigger words drawn from a small vocabulary (500–1000 tokens). The target label is the absolute index distance between these words. This task demands tight coupling of token identity and relative position. We generate 100,000 examples with a 7:3 train/test split.
Task 2: Trigger Word Counting. Each sequence is labeled by the total count of occurrences of a fixed set of 10–30 trigger words, irrespective of their positions. This position-agnostic task tests pure content recognition. We generate 50,000 examples with a 7:3 train/test split.
These tasks enable a controlled comparison of PE methods under position-sensitive (Task 1) and position-agnostic (Task 2) settings, directly testing our theoretical framework.
4.2 Convergence Analysis
We plot the experimental results as line graphs. The training accuracy curve indicates the convergence speed of each PE method on the training set, while the test accuracy curve reflects its capacity to learn the task’s underlying rules. Convergence results for Task 1 and Task 2 are displayed in Figures 2 and 3, respectively.
Figure 2: Training and test accuracy vs. epoch for different PE methods on Task 1 (relative-dis).
Task 1: Trigger Word Relative Distance Prediction. Leveraging its Hadamard-product coupling, RoPE converges fastest and most stably, reaching high training and test accuracy with a minimal generalization gap. By contrast, Absolute encoding ranks second: its fixed first-layer bias $G_{p}$ can be adjusted dynamically in deeper layers. Random initialization matches Absolute on training accuracy but overfits—its non-Toeplitz bias limits generalization, yielding lower test performance. Meanwhile, ALiBi performs poorly: its fixed, data-agnostic Toeplitz bias $B$ impedes adaptation and slows convergence. Neither NoPE nor standard Relative generalizes, since both lack a mechanism to bind content and relative position.
Figure 3: Training and test accuracy vs. epoch for different PE methods on Task 2 (counting).
Task 2: Trigger Word Count. In the position-agnostic Task 2, imposing positional biases generally degrades generalization. Absolute encoding incurs a mild penalty—its fixed $G_{p}$ can still be refined by later layers—while RoPE’s soft, data-adaptive bias imposes even less overhead, yielding superior test accuracy. Overfitting emerges under Random initialization: its non-Toeplitz bias perfectly fits content counts but produces the largest train–test gap. Relative encoding attenuates its learned bias toward zero and thus matches NoPE’s performance. ALiBi suffers most, as its rigid, fixed Toeplitz bias obstructs even simple counting.
@@ 5 Ablation Studies and Mechanistic Insights
5.1 Attention Head Specialization for Positional Processing
To complete our theoretical argument, we predict that PE mechanisms with spectrally stable position coupling—such as RoPE—will localize essential positional computations into a few specialized attention heads, especially in early layers, yielding a deposit pattern. We test this via a causal, head-wise ablation study: for each layer $l$ and head $h$, we zero out the head’s output $H_{l,h}$ and measure the resulting drop in Task 1 test accuracy relative to the intact model. This accuracy reduction quantifies each head’s criticality for content–position coupling.
Figure 4: Head-wise ablation violin plot for three situations.
Figure 4 confirms that under RoPE, ablating a single shallow-layer head causes a pronounced accuracy loss, indicating early-layer specialization. By contrast, no pronounced deposit pattern appears under No-RoPE methods like NoPE or ALiBi. In the main text we illustrate NoPE’s Task 1 ablation and RoPE’s Task 2 ablation in Figure 4; the full set of non-RoPE results (Absolute, ALiBi, Relative, Random) is provided in Appendix C.
Building on this finding, we hypothesize that only a subset of heads needs RoPE to capture relative position, while the remaining heads can operate without positional bias. We therefore conduct a “partial RoPE” experiment, applying RoPE to $k$ heads and NoPE to the rest. As Figure 5 shows, even with only a handful of RoPE-enabled heads, Task 1 performance remains below-near-optimal, and Task 2 generalization is shocked.
@@ 5.1 Attention Head Specialization for Positional Processing
Figure 5: Performance when RoPE is applied to a subset of heads.
5.2 Modulating Deposit Patterns via MLA
Motivated by the observed deposit pattern and the “partial RoPE” results, we explore two strategies to diffuse positional responsibility across heads. First, we augment RoPE by adding an absolute positional embedding before the first layer, which provides an initial position cue that can be adjusted by deeper layers. Second, we adopt the MLA (Multi-head Latent Attention) mechanism from Deepseek-V3, which mixes NoPE and RoPE signals before the attention inner product.
We present the result of the early method in Figure 6. Due to the absolute position encoding, the deposit pattern no longer appears in the first layer. Ablating each head does not affect the overall performance. However, starting from the second layer, the deposit pattern reappears, which also indirectly reflects that the deposit pattern is an inherent phenomenon of the coupling between rope capture content and relative position. It will not be completely eliminated by adding absolute position information to the input.
Figure 6: (Left) The performance of RoPE with an absolute position on Task 1 and Task 2. (Right) "Deposit Pattern" of RoPE-Abs.
The late method MLA could solve the deposit pattern. Firstly, we use the inner integral calculation of an attention head to demonstrate MLA:
$<q_{i},k_{j}>=<(W_{UQ}W_{DQ}x_{i};R_{i}W_{UR}W_{DQ}x_{i}),(W_{UK}W_{DKV}x_{j};R_{j}W_{KR}x_{j})>,$ (11)
where (;)(\;;\;)( ; ) means the operation of vector stitching. Without considering the dimensionality reduction operation of MLA to improve computational efficiency, we can understand that each head independently processes the inner product of NoPE and RoPE once and sums them up, and the parameters of the two parts are independent and do not interfere with each other. In our theoretical frame work, it can be expressed as
@@ 5.2 Modulating Deposit Patterns via MLA
$\mathbf{L}_{\text{MLA\_complex}}=\left(G_{\mathbf{q},\mathbf{k}}+G_{\mathbf{q},\mathbf{p}}+G_{\mathbf{p},\mathbf{k}}+G_{\mathbf{p}}\right)\circ(\alpha G_{\mathbf{e}}+I).$ (12)
We transplant this MLA module into our 6-layer decoder (see Appendix C for hyperparameters). Table 1 shows the MLA performance on the two tasks. Crucially, head-wise ablation under MLA (Figure 7) shows no single head is indispensable for Task 1, demonstrating effective mitigation of the earlier deposit pattern.
Figure 7: Head-wise ablation violin plot under MLA.
@@
Although RoPE was designed to enhance length generalization, real-world results often fall short of this ideal. We hypothesize that early single-head deposition hinders length robustness: when one head monopolizes positional processing, its representations may not effectively transfer to deeper layers under length shifts. The MLA mechanism, by diffusing positional responsibility, points toward a remedy. More broadly, our framework suggests that transforming the relative-position signal $G_{e}$ into a distributed form $F(G_{e})$—for example, via controlled mixing—could reconcile spectral stability with robust, length-agnostic encoding. We provide a specific qualitative statement of this view in Appendix D.

### v2

@@ 3.2 A Unified Framework via Toeplitz Matrices
The complex form of RoPE is an important component of the original paper (Su et al., 2024). This form, combined with the Abel summation, can prove an upper bound on its distance decay. A detailed derivation of how to convert the complex RoPE form into the Toeplitz matrix form above is provided in Appendix B. This formulation highlights that RoPE does not add position information; instead, it modulates all content interactions through a shared position-dependent kernel.
This formal analysis reveals a critical distinction. Additive methods combine content and position signals, but RoPE’s multiplicative structure uses a Toeplitz kernel ($G_{\mathbf{e}}$) to modulate the entire content-content and content-position interaction. This enables a more direct and expressive coupling, allowing the model to learn attention patterns where the relevance of a token’s content is conditioned on its relative position to another. This type of task cannot be achieved through pure translation invariance; instead, it requires that a specific translation distance be directly affected by the content feedback. We hypothesize that this powerful coupling mechanism comes with a drawback: by routing the learning of positional dependencies through this multiplicative kernel, RoPE may create a strong inductive bias that encourages positional specialization, concentrating this logic into a small subset of attention heads. This hypothesis motivates the experiments in the following section.
4 Experiments
To empirically test our hypothesis about the distinct behaviors of additive and multiplicative positional encodings, we design two synthetic tasks. These tasks allow us to isolate the model’s ability to couple token content with positional information in a controlled setting. We use a 6-layer Transformer decoder as our base model and evaluate six PE configurations: Absolute (Vaswani et al., 2017), T5 Relative Bias (Raffel et al., 2020), ALiBi (Press et al., 2022), RoPE (Su et al., 2024), NoPE (no explicit PE) (Kazemnejad et al., 2023), and a Random initialized embedding baseline. Our experiments are averaged over five seeds, and the variance is less than 5%. Full implementation and training details are provided in Appendix D.
4.1 Synthetic Task Design
Task 1 – Position-Sensitive: Relative Distance Classification. This task is designed to require strong content-position coupling, the theorized strength of multiplicative mechanisms. Each input sequence contains two unique "trigger" words drawn from a vocabulary of 1,000 tokens. The model must predict the relative distance between them, formulated as a classification problem over binned distances. Success requires the model to identify what the trigger words are and determine where they are in relation to each other. We generated a dataset of 100,000 examples for this task.
Task 2 – Position-Agnostic: Trigger Word Counting. This task serves as a control, designed to penalize any inflexible positional bias. The model must predict the total number of occurrences of a predefined set of 20 trigger words within a sequence, regardless of their positions. Because positional information is a nuisance variable, an ideal model should learn to ignore it. We generated a dataset of 50,000 examples for this task.
4.2 Performance and Generalization Analysis
We first analyze the learning dynamics and generalization performance of each PE method on our two tasks. The results are presented in Figures 2 and 3.
Figure 2: Training and test accuracy on Task 1 (Position-Sensitive). RoPE’s multiplicative coupling enables it to significantly outperform all other methods in both convergence speed and final generalization.
On the Position-Sensitive Task, RoPE Excels. As predicted by our framework, RoPE (Figure 2) demonstrates superior performance. It converges fastest and achieves the highest test accuracy with a minimal generalization gap. This supports our theory that its multiplicative content-position coupling is uniquely suited for tasks demanding this form of reasoning. In contrast, additive methods show mixed results. Absolute PE performs second best, as its initial positional signal can be adapted by subsequent layers. The Random embedding baseline learns the training set but fails to generalize, as its non-Toeplitz structure lacks the translation invariance required to learn a general rule. Methods like ALiBi, with its fixed and data-agnostic Toeplitz bias, struggle to adapt, while NoPE and standard Relative PE fail entirely, lacking any mechanism to bind content to position.
Figure 3: Training and test accuracy on Task 2 (Position-Agnostic). Strong, inflexible positional biases like ALiBi are detrimental. RoPE’s adaptive nature allows it to perform well, nearly matching methods with no positional bias.
On the Position-Agnostic Task, Strong Positional Biases are Detrimental. Figure 3 confirms our second prediction: on a task where position is irrelevant, a strong positional inductive bias is harmful. ALiBi, with its rigid, non-adaptive bias, suffers the most severe performance degradation. Conversely, NoPE and Relative PE (which learns to attenuate its bias towards zero) perform well, as they do not impose a counterproductive positional structure. RoPE also performs strongly, demonstrating its adaptability; its multiplicative mechanism allows the model to learn to down-weight the positional signal when it is not useful. Absolute PE incurs a small penalty, while the Random baseline again overfits, learning spurious correlations between positions and counts in the training data.
4.3 Evidence of the Single-Head Deposit Pattern in RoPE
Our theory suggests that RoPE’s multiplicative structure concentrates positional reasoning into a few specialized heads. To find direct evidence for this phenomenon, we conduct a head-wise causal ablation study on the models trained on Task 1. For each head in the network, we zero out its output and measure the resulting drop in test accuracy. A large drop indicates that the head is critical to the task.
Figure 4: Head-wise ablation reveals a unique "Deposit Pattern" in RoPE. Each violin shows the distribution of accuracies after ablating each head in that layer; long vertical tails correspond to outlier heads whose removal causes a large accuracy drop. The three insets illustrate the characteristic shapes of these distributions: a “deposit” shape with a single deep outlier (RoPE–Task 1), a flat “average” shape (NoPE–Task 1), and a noisy “shock” shape (RoPE–Task 2). These shapes match the contours of the main violins. Full per-head curves are provided in Appendix D.2.
The results, shown in Figure 4, provide striking confirmation of our hypothesis. In the RoPE model, nearly all positional logic is localized to a single head in the first layer. Ablating this one head causes a catastrophic drop in accuracy ($\approx$60%), while ablating other heads has a negligible effect. We term this phenomenon the “single-head deposit pattern.” Crucially, this pattern is unique to the combination of the RoPE architecture and a position-sensitive task. As shown in Figure 4, the pattern does not emerge in the NoPE model on the same task, nor does it appear in the RoPE model trained on the position-agnostic Task 2. This confirms that the deposit pattern is not a generic artifact of Transformer training but a direct consequence of RoPE’s multiplicative content-position coupling mechanism. We therefore interpret the deposit pattern as a training bias rather than clean modularity, although a full causal link to length generalization is left for future work (Appendix E).
@@ 5 Ablation Studies and Mechanistic Insights
The discovery of the "single-head deposit pattern" raises critical questions about the underlying mechanics of RoPE. Is this pattern an inevitable side effect of multiplicative coupling, an efficient specialization, or a wasteful bottleneck? To elucidate the mechanisms driving this phenomenon, we design three targeted ablation studies, each testing a specific hypothesis.
5.1 Ablation 1: Does RoPE Suppress Latent Positional Representations?
Hypothesis: We hypothesize that RoPE’s explicit, multiplicative positional signal is so potent that it inhibits the model from forming its own implicit positional representations—a known capability of standard Transformers (Zuo et al., 2025; Kazemnejad et al., 2023)—thereby causing other heads to remain position-agnostic.
Experimental Design: To test this, we inject a redundant signal by augmenting a RoPE-equipped Transformer with an additional Absolute PE at the input layer. This directly provides an explicit $p_{i}$ component before the RoPE-enabled attention layers. We then observe if this alters the deposit pattern.
Figure 5: (Left) Performance of a hybrid RoPE + Absolute PE model. (Right) The deposit pattern is initially suppressed in Layer 1 but re-emerges in deeper layers, demonstrating the strong inductive bias of RoPE.
Results: The results, shown in Figure 5, are revealing. The explicit Absolute PE signal disrupts the deposit pattern in the first layer, distributing positional responsibility more evenly. However, the pattern re-emerges in deeper layers. This strongly suggests that while an initial explicit signal can be utilized, the powerful inductive bias of RoPE’s multiplicative coupling eventually reasserts itself, overriding the additive signal. This confirms our hypothesis: RoPE’s mechanism is so dominant that it suppresses the model’s tendency to form or utilize other forms of positional representation.
5.2 Ablation 2: Is a Single RoPE-Enabled Head Sufficient?
Hypothesis: The deposit pattern implies that for tasks like ours, a single RoPE-enabled head is sufficient for positional reasoning, rendering additional RoPE heads redundant.
Experimental Design: We test this by systematically reducing the number of attention heads that utilize RoPE, from all heads down to just one, with the rest operating as NoPE heads. We evaluate performance on Task 1.
@@ 5.2 Ablation 2: Is a Single RoPE-Enabled Head Sufficient?
Figure 6: Performance on Task 1 when RoPE is applied to a diminishing subset of heads. Performance remains near-optimal even with just one or two RoPE-enabled heads.
Results: As shown in Figure 6, performance on Task 1 remains near-optimal even when only one or two heads are equipped with RoPE. This confirms that a small number of specialized heads are indeed sufficient to capture the necessary positional information for this task. The result supports our explanation for the deposit pattern: once one head effectively learns the content-position interaction via the multiplicative mechanism, other heads become inert with respect to this task, as there is no learning incentive to develop redundant positional capabilities.
5.3 Ablation 3: Can a Hybrid Architecture Mitigate Inefficiency?
Hypothesis: The deposit pattern, while effective, is an inefficient use of model capacity. We hypothesize that a hybrid architecture can mitigate this by explicitly dedicating resources, retaining RoPE’s strengths while improving robustness.
Experimental Design: To explore this, we evaluate the Multi-Latent Attention (MLA) architecture from DeepSeek (Liu et al., 2024). MLA formalizes a hybrid approach by creating parallel pathways for position-agnostic and position-sensitive processing within a single head. Formally, the query-key inner product is structured as a concatenation of a NoPE component and a RoPE component:
$\displaystyle\langle q_{i},k_{j}\rangle=\langle(W_{UQ}x_{i}\,;\,R_{i}W_{UR}x_{i}),(W_{UK}x_{j}\,;\,R_{j}W_{UR}x_{j})\rangle,$ (3)
where $(\,;\,)$ denotes vector concatenation and the projection matrices for each pathway are distinct. This structure ensures that each attention head computes both a NoPE-style and a RoPE-style logit simultaneously. Interpreted through our theoretical framework, this mechanism effectively modifies the logit matrix by creating a weighted combination of a standard content matrix and a RoPE-modulated one:
@@ 5.3 Ablation 3: Can a Hybrid Architecture Mitigate Inefficiency?
$\displaystyle\mathbf{L}_{\text{MLA\_complex}}=\left(G_{\mathbf{q},\mathbf{k}}+G_{\mathbf{q},\mathbf{p}}+G_{\mathbf{p},\mathbf{k}}+G_{\mathbf{p}}\right)\circ(\alpha G_{\mathbf{e}}+I).$ (4)
This prevents the purely multiplicative signal from dominating and encourages a more distributed representation of position. We replaced our standard attention with the MLA module and evaluated it on both tasks.
Figure 7: Head‐wise ablation violin plot under MLA.
@@
Results: The MLA architecture successfully mitigates the deposit pattern. Figure 5.3 shows that no single head is indispensable, and positional responsibility is diffused across the model. Furthermore, Table 5.3 shows that MLA nearly matches RoPE’s performance on Task 1 while dramatically improving on Task 2. This demonstrates that by preventing over-specialization, the hybrid model achieves a more robust balance.
5.4 Ablation 4: Where Do Additive and Multiplicative PEs Store Positional Information?
Hypothesis. Given the decomposition $x_{i}=c_{i}+p_{i}$ (Assumption 3.1), we ask whether different PE mechanisms retain an explicit positional direction $p_{i}$ in the activations, or whether this information is absorbed into the query–key parameterization. Our hypothesis is that additive PEs preserve activation-level $p_{i}$, while RoPE rapidly suppresses it.
Experimental Design. For each layer $\ell$, we ablate either (1) the fixed positional vectors $p_{i}$ (Absolute PE) or (2) norm-matched random vectors, and measure test accuracy. A larger drop under $p_{i}$-removal indicates that the model relies on an explicit positional direction. For RoPE, we repeat the same intervention on the hybrid Absolute+RoPE model (Section 5.1), where $p_{i}$ is added only at the embedding layer.
Figure 8: Layer-wise ablation of absolute positional vectors versus norm-matched random vectors. Left: Absolute PE shows consistently larger drops under $p_{i}$-removal, confirming that additive PEs preserve explicit positional directions in the activations. Right: For Absolute+RoPE, this difference vanishes after Layer 2, indicating that RoPE suppresses activation-level positional directions and shifts positional structure into the attention parameterization.
Results. Fig. 8 shows a clear contrast: (1) For Absolute PE, removing $p_{i}$ causes a consistently larger accuracy drop than removing norm-matched random vectors, indicating that additive PEs preserve explicit positional directions in the activations; (2) For Absolute+RoPE, this difference disappears after Layer 2, showing that RoPE suppresses activation-level positional directions and shifts positional information into the query–key parameterization.
Summary. Additive PEs store position in explicit activation directions $p_{i}$, whereas RoPE rapidly eliminates these directions and embeds positional structure in the frequency-modulated parameters. This is why the structure of MLA can eliminate this phenomenon.
6 Theoretical Analysis of the Deposit Pattern
The empirical discovery of the "single-head deposit pattern" in RoPE (Su et al., 2024) invites a formal theoretical explanation. In this section, we provide a mathematical argument demonstrating that this phenomenon is an inherent consequence of RoPE’s design. We analyze the gradient signal for the relative distance classification task (Task 1). Let $(i_{s},j_{s})$ be the positions of the trigger words for a sample $s$, with the task being to predict the relative distance $d_{s}:=|j_{s}-i_{s}|$. The algebraic intuition behind our proof is this: for a sample labeled $|i-j|$, the optimization rewards the distance signal $|i-j|$ and penalizes other distance signals. For addition, the reward and penalty are always antagonistic because the position bias lacks context information. However, for multiplication, the penalty is attenuated by frequency, making the reward cumulative. This accumulation ultimately leads to an initial top-level bias in the gradient signal, which is amplified by backpropagation. As a comparison with RoPE, we also analyze ALiBi (Press et al., 2022) on this task.
Our analysis relies on the following idealized assumptions, which are empirically plausible in the context of early training.
Assumption 6.1 (Formal Setup for Gradient Analysis).
  1. (A1)
    Jacobian Stability. Each inter-layer Jacobian $J^{(l)}$ has singular values bounded in $[1-\alpha_{l},1+\alpha_{l}]$ for $\alpha_{l}\in(0,1)$, consistent with architectures using pre-normalization.
  2. (A2)
    Monotone Anchor Path. For each sample $s$, there is an "anchor" row $u_{s}$ where the gradient signal is positively correlated with the target token’s value vector. Formally, $\langle\partial\ell_{s}/\partial Y^{(h,L)}_{u_{s},:},V^{(h,L)}_{j_{s},:}\rangle\geq\eta_{s}^{(h)}>0$. This formalizes that a basic learning signal exists.
  3. (A3)
    Projection Separability (for ALiBi). The projection matrices $W_{Q},W_{K},W_{V}$ are such that $\ker W_{Q}\cap\ker W_{K}\not\subset\ker W_{V}$. This is a generic condition in overparameterized models.
Our analysis demonstrates a fundamental difference in how RoPE and ALiBi structure the gradient signal for relative position tasks.
Proposition 6.1 (RoPE Gradient Seed Coherence).
Under Assumption (A2), the aggregated anchor gradient for RoPE, $H^{(h)}_{L}(d)$, has a deterministic lower bound. A strictly positive seed ($H^{(h)}_{L}(d)>0$) is guaranteed if the learning signal is sufficiently strong to satisfy the condition $a_{*}^{(h)}(d)\,\eta_{*}^{(h)}(d)>C^{(h)}\chi_{L}$.
Proposition 6.2 (ALiBi Gradient Cancellation).
Under Assumptions (A3), for any empirical distance distribution $\hat{\mu}_{N}$, there exists a batch of samples realizing $\hat{\mu}_{N}$ such that the aggregated anchor gradient for ALiBi vanishes for all heads and distances: $H^{(h)}_{L}(d)\equiv 0$.
The existence of a non-cancellable seed in RoPE is the starting point. The following theorem shows how this seed inevitably leads to the deposit pattern.
Theorem 6.1 (Exponential Amplification and Specialization).
Let $\mathcal{U}$ be the subspace of gradient directions corresponding to the positive RoPE seeds from Proposition 6.1. The signal-to-noise ratio of the gradient within this subspace is amplified exponentially during backpropagation. Crucially, the margin between the gradient signal for the dominant head and that of the second-dominant head also grows exponentially.
$\text{Margin}_{l}\geq\text{Margin}_{L}\prod_{k=l}^{L-1}\gamma_{k},\quad\text{where each gain factor }\gamma_{k}>1.$ (5)
Implication.
Theorem 6.1 provides the theoretical basis for the single-head deposit pattern. Even a minuscule initial advantage for one head in learning the position-sensitive task (a non-zero MarginL) becomes exponentially magnified as gradients propagate to lower layers. This dynamic strongly incentivizes the optimizer to allocate the entire task of positional reasoning to that single head, as its parameters receive a disproportionately larger learning signal, thus explaining the observed specialization. The detailed proofs for these results are provided in Appendix C.